OpenStudy (anonymous):

A pair of jeans is 32 1/2 inches long. How long are the jeans after they have been washed if they shrink 1 1/3 inches?

OpenStudy (anonymous):

that would be ( 32 1/2 - 1 1/3 = 32 - 1 = 31) ...... so 31 1/2 - 1/3... 31 1/6

OpenStudy (anonymous):

Thank you for your help

OpenStudy (anonymous):

np